OpenStudy (anonymous):

what is 2x^2 -7 +3=0??

OpenStudy (anonymous):

2x^2 -7 +3=0 2x^2 = 7 -3 2x^2 = 4 x^2 = 4/2 x = sqrt(2)

OpenStudy (anonymous):

2(x^2-2) x= + or - sqrt(2)

OpenStudy (mayankdevnani):

2x^2-4=0 2x^2=4 x^2=4/2=2 x=under root2

OpenStudy (mayankdevnani):

ok @MusicTeacher

OpenStudy (mayankdevnani):

answer is \[\sqrt{2}\]

OpenStudy (anonymous):

ok ty but the answer is supposed to be 1/2

OpenStudy (mayankdevnani):

\[x^2-7+3=0\] is it your question?

OpenStudy (mayankdevnani):

@MusicTeacher

OpenStudy (anonymous):

no its \[\pm \sqrt{2}\]